# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v3

@@ 5 Experiments
We compare DDVI with Auto-Encoding Variational Bayes (AEVB) [14], AEVB with inverse autoregressive flow posteriors (AEVB-IAF) [16], Adversarial Auto-Encoding Bayes (AAEB) [20], and Path Integral Sampler (PIS) [46] on MNIST [18] and CIFAR-10 [17] in unsupervised and semi-supervised learning settings, and also on the Thousand Genomes dataset [35]. We also compare with Hierachical Auto-Encoding Variational Bayes (H-AEVB) [26, 38] in unsupervised setting. We discuss the computational costs of all methods in Appendix D. The priors, model architecture, and training details can be founded in Appendix F, Appendix G, and Appendix H respectively. All results below are reported with 95% confidence interval using 3 different seeds.
@@ 5.1 Unsupervised learning
We start with synthetic experiments that are aimed at benchmarking the expressivity of diffusion-based posteriors and their ability to improve fitting $p$, a distribution with a complex structured prior, like one might find in probabilistic programming, scientific analysis, or other applications. We fit a model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and Table 6. The model distribution $p_{\bm{\theta}}$ is instantiated by a deep Gaussian latent variable model (DGLVM) with multi-layer perceptrons (MLPs) on MNIST and convolutional neural networks (CNNs) on CIFAR-10. The details of model architecture are provided in Appendix G.
Our first set of metrics (ELBO and MMD) seeks to evaluate the learned generative model $p_{\bm{\theta}}$ is good. In the ELBO calculation, we average the reconstruction loss across image pixels. We use MMD to measure sample quality: we generate images with the trained model and calculate MMD between the generated images and test images using a mixture of Gaussian kernel111with sigma equal to [2, 5, 10, 20, 40, 80]. We only report MMD for MNIST, since CIFAR-10 generated samples are very low-quality for all methods because the latent dimension is 2.
@@
From Tables 1 and 6, we see our method DDVI achieve best ELBO in all but one scenario, in which it still performs competitively. We also see strong results in Latent NLL and Acc in many scenarios, except for Swiss Roll where AAEB does well. We present visualizations of MNIST using the baseline methods and our method in Figure 2.
@@
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-11.15\pm 0.53$ $0.93\pm 0.01$ $1.36\pm 0.03$ $-15.29\pm 1.33$ $0.68\pm 0.01$ $4.60\pm 0.23$ $-10.26\pm 0.25$ $0.86\pm 0.01$ $1.68\pm 0.02$ AEVB-IAF $-2.10\pm 0.26$ $\bm{0.95\pm 0.00}$ $\bm{1.06\pm 0.03}$ $-5.38\pm 1.78$ $0.90\pm 0.02$ $2.75\pm 0.14$ $-2.67\pm 0.83$ $\bm{0.91\pm 0.01}$ $\bm{0.90\pm 0.02}$ AAEB $N/A$ $0.89\pm 0.01$ $1.55\pm 0.01$ $N/A$ $0.88\pm 0.01$ $3.07\pm 0.05$ $N/A$ $1.94\pm 0.38$ $0.76\pm 0.13$ DDVI $\bm{-0.24\pm 0.13}$ $\bm{0.95\pm 0.00}$ $\bm{1.06\pm 0.01}$ $\bm{-2.89\pm 0.33}$ $\bm{0.92\pm 0.01}$ $\bm{2.09\pm 0.00}$ $\bm{0.02\pm 0.09}$ $0.90\pm 0.01$ $1.49\pm 0.03$
@@ 5.2 Semi-supervised Learning
For this setting, we evaluate ELBO, latent negative log-likelihood (Latent NLL), and a k-nearest neighbors classification accuracy of the latents (Acc). We choose classification accuracy since classification is a common downstream task for semi-supervised learning. We use the same set of priors and baselines. Details on how we partition each prior into $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|\mathbf{x},l)$ can be founded in Appendix F. The partitions defined for our priors are local parts of the priors. We note that unlike unsupervised learning, we use the simplified sleep term from Appendix B in our objective for this setting, since $q_{\bm{\phi}}$ already gets extra information from $l$ here.
The results are shown in Tables 2 and 7. DDVI mostly outperforms the baselines across different priors and metrics, especially on CIFAR-10 where DDVI is best across the board. For MNIST, DDVI always achieves the best ELBO, and it also performs competitively with other baselines in classification accuracy. We also show the visualizations of the latents in Figure 3 where DDVI matches the prior almost perfectly.
@@ 5.3 Clustering and Visualization for Genotype Analysis
Method Cluster Cluster NMI
Method Purity Completeness
AEVB $0.28\pm 0.02$ $\bm{0.78\pm 0.16}$ $0.59\pm 0.08$
AEVB-IAF $0.29\pm 0.04$ $\bm{0.73\pm 0.06}$ $0.55\pm 0.06$
AAEB $0.37\pm 0.06$ $\bm{0.76\pm 0.11}$ ${0.63\pm 0.02}$
DDVI $\bm{0.45\pm 0.03}$ $\bm{0.75\pm 0.05}$ $\bm{0.66\pm 0.04}$
Figure 5: Quantitative genotype clustering results.
In Figure 5, we see that DDVI attains the best performance on cluster purity and NMI. For cluster completeness, VAE and AAE have better means but much larger confidence interval. Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA [44], t-SNE [40], and UMAP [23]). The legend of Figure 4 can be founded at Figure 6.
@@ 6 Discussion
Our approach is most similar to flow-based approximators [28, 16]; in fact when $T\to\infty$, our diffusion-based posterior effectively becomes a continuous-time normalizing flow [36]. However, classical flow-based methods require invertible architectures for each flow layer: this constrains their expressivity and requires backpropagating through potentially a very deep network. Our approach, on the other hand, trains a model (a continuous-time flow when $T\to\infty$) via a denoising objective (similar to score matching) that does not require invertible architectures and effectively admits an infinite number of layers (with weight sharing). This model is trained not by backpropagating through the ELBO, but rather via an auxiliary diffusion loss term (effectively, a score matching objective).
@@
Variational posteriors based on GANs [20] also admit expressive architectures and require only sample-based access to the prior $p(\mathbf{z})$. Our diffusion-based approach admits a more stable loss, and is potentially more expressive, as it effectively supports an infnite number of layers (with shared parameters when $T\to\infty$). Unlike GANs, our models also admit explicit likelihoods and allow us to compute the ELBO for model evaluation. Our approach is similar to variational MCMC [31]; however, we train with a better objective augmented with a diffusion loss, and we adopt improved architectures with shared weights across layers.
@@
Our algorithms are also simpler. For example, diffusion sampling methods require backpropagating through a sampling process to minimize the reverse $\text{KL}(q_{\bm{\phi}}||p_{\bm{\theta}})$, which poses challenges with optimization and credit assignment. Some methods based on Schrodinger bridges require an iterative optimization process generalizing the sinkhorn algorithm or computationally expensive on-policy or off-policy [21] trajectory-based optimization. In contrast, DDVI optimizes the forward $\text{KL}(p_{\bm{\theta}}||q_{\bm{\phi}})$ using simple gradient-based optimization that directly emulates diffusion-based training.
@@ 7 Related Work
Vahdat et al. [39], Wehenkel and Louppe [43], Rombach et al. [30] perform diffusion in the latent space of a VAE to improve the efficiency of image generation. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors. Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
Recent work [25, 47, 42] has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. Cohen et al. [5] crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
@@
Diffusion sampling [3, 46, 41, 45, 29, 32, 1] mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) [46], denoising diffusion sampler (DDS) [41], and time-reversed diffusion sampler (DIS) [3], which have been unified by Richter et al. [29]. Some other works [45, 1] use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. Sendera et al. [32] benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps.
@@
Latent variable models in general are an attractive alternative to visualization methods like PCA, UMAP, and t-SNE [23, 40]. Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space. Nevertheless, downsides of LVMs are that they are more computationally expensive and require careful hyperparameter tuning.
@@ 8 Conclusion
While this paper focuses on applications of DDVI to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality. Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making [24, 6], meta-learning [27], or causal effect estimation [7]. Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models. Also, attaining competitive likelihood estimation requires architecture improvements that are orthogonal to this paper. However, our ability to generate diverse samples and achieve class separation in latent space hints at the method’s potential on these tasks.
@@ Appendix A Pseudocode
Here we provide a pseudocode to illustrate the training process of DDVI. We release anonymized code repo at https://anonymous.4open.science/r/vi-with-diffusion-D70F.
@@
1:(Optional) Pre-train $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$ and $q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})$ with DDVI but with unconditional diffusion model $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$
@@
4:         $\mathbf{y}_{i}\sim q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}_{i})$ and $\mathbf{z}_{i}\sim q(\mathbf{z}|\mathbf{y}_{i},\mathbf{x}_{i})$ for $i=1,\ldots,k$
5:         Optimize ${\bm{\theta}},{\bm{\phi}}$ with respect to a Monte Carlo estimate of $\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\left(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\|p_{\bm{\theta}}(\mathbf{y},\mathbf{z})\right)$ for each $\mathbf{x}_{i}$triangleright Standard ELBO training part
@@
10:              Optimize ${\bm{\phi}}$ using the standard diffusion noise prediction loss on $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{i},\hat{\mathbf{x}}_{i})$
@@ Appendix B Simplifying Wake-Sleep
In wake-sleep, sampling $\mathbf{x}$ from $p_{\bm{\theta}}$ to obtain gradients for the sleep term introduces computational overhead. To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ of $\mathcal{L}$:
@@
^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$}}.$ (16)
@@
We have replaced $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ with a latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$, in which $\mathbf{x}$ is given, and we only seek to fit the true reverse noising process $r(\mathbf{z}|\mathbf{y})$ independently of $\mathbf{x}$. We can similarly show that
$\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$ $=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})r(\mathbf{y}|\mathbf{z})}[\log({q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}/r(\mathbf{y}|\mathbf{z}))]+\bar{H}(p_{\bm{\theta}})$ (17)
$=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[D_{\mathrm{KL}}(r(\mathbf{y}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q(\mathbf{z}|\mathbf{x})),$ (18)
@@
where $\bar{H}(p_{\bm{\theta}})$ is an entropy term constant in ${\bm{\phi}}$. Thus, we minimize the forward KL divergence by sampling $\mathbf{z}$, and applying the noising process to get $\mathbf{y}$; the $q_{\bm{\phi}}$ is fit to denoise $\mathbf{z}$ from $\mathbf{y}$ as in Equation 10.
@@
We optimize our bound on ^$\hat{\mathcal{L}}(x,{\bm{\theta}},{\bm{\phi}})$ end-to-end using minibatch gradient descent over ${\bm{\theta}},{\bm{\phi}}$. While the wake term is a reconstruction loss as in wake-sleep, the sleep term generates latent samples $\mathbf{z},\mathbf{y}$ from $r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ (by analogy with $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ in normal wake-sleep); the denoiser $q_{\bm{\phi}}$ is trained to recover $\mathbf{z}$ from $\mathbf{y}$. Thus, we perform wake-sleep in latent space, which obviates the need for alternating wake and sleep phases, and allows efficient end-to-end training. A limitation of this approximation is that the sleep term does not fit $q_{\bm{\phi}}$ to the true $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x},\mathbf{y})$, and as a consequence $\hat{L}$ is not a tight lower bound on $\log p_{\bm{\theta}}(\mathbf{x})$. We may think of $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$ as a regularizer to the ELBO.

### v4

@@ 5 Experiments
We compare DDVI with Auto-Encoding Variational Bayes (AEVB) (Kingma and Welling 2013), AEVB with inverse autoregressive flow posteriors (AEVB-IAF) (Kingma et al. 2016), Adversarial Auto-Encoding Bayes (AAEB) (Makhzani et al. 2015), and Path Integral Sampler (PIS) (Zhang and Chen 2021) on MNIST (Lecun et al. 1998) and CIFAR-10 (Krizhevsky and Hinton 2009) in unsupervised and semi-supervised learning settings, and also on the Thousand Genomes dataset (Siva 2008). We also compare with Hierachical Auto-Encoding Variational Bayes (H-AEVB) (Ranganath, Tran, and Blei 2016; Vahdat and Kautz 2020) in unsupervised setting. We discuss the computational costs of all methods in Appendix D. The priors, model architecture, and training details can also be founded in Appendix H. All results below are reported with 95% confidence interval using 3 different seeds.
@@ 5.1 Unsupervised learning
We start with synthetic experiments that are aimed at benchmarking the expressivity of diffusion-based posteriors and their ability to improve fitting $p$, a distribution with a complex structured prior, like one might find in probabilistic programming, scientific analysis, or other applications. We fit a model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and 7. The model distribution $p_{\boldsymbol{\theta}}$ is instantiated by a deep Gaussian latent variable model (DGLVM) with multi-layer perceptrons (MLPs) on MNIST and convolutional neural networks (CNNs) on CIFAR-10 (see the details of model architecture in Appendix G).
Our first set of metrics (ELBO and MMD) seeks to evaluate the learned generative model $p_{\boldsymbol{\theta}}$ is good. In the ELBO calculation, we average the reconstruction loss across image pixels. We use MMD to measure sample quality: we generate images with the trained model and calculate MMD between the generated images and test images using a mixture of Gaussian kernel with sigma equal to $[2,5,10,20,40,80]$. We only report MMD for MNIST, since CIFAR-10 generated samples are very low-quality for all methods because the latent dimension is 2.
@@
From Table 1 and Table 7 in Appendix, we see our method DDVI achieve best ELBO in all but one scenario, in which it still performs competitively. We also see strong results in Latent NLL and k-nearest neighbors classification accuracy of the latents (Acc). in many scenarios, except for swiss roll where AAEB does well. We present visualizations on MNIST using the baselines and our method in Figure 2.
@@
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-11.15\pm 0.53$ $0.93\pm 0.01$ $1.36\pm 0.03$ $-15.29\pm 1.33$ $0.68\pm 0.01$ $4.60\pm 0.23$ $-10.26\pm 0.25$ $0.86\pm 0.01$ $1.68\pm 0.02$ AEVB-IAF $-2.10\pm 0.26$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.03}$ $-5.38\pm 1.78$ $0.90\pm 0.02$ $2.75\pm 0.14$ $-2.67\pm 0.83$ $\mathbf{0.91\pm 0.01}$ $\mathbf{0.90\pm 0.02}$ AAEB $N/A$ $0.89\pm 0.01$ $1.55\pm 0.01$ $N/A$ $0.88\pm 0.01$ $3.07\pm 0.05$ $N/A$ $1.94\pm 0.38$ $0.76\pm 0.13$ DDVI $\mathbf{-0.24\pm 0.13}$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.01}$ $\mathbf{-2.89\pm 0.33}$ $\mathbf{0.92\pm 0.01}$ $\mathbf{2.09\pm 0.00}$ $\mathbf{0.02\pm 0.09}$ $0.90\pm 0.01$ $1.49\pm 0.03$
@@ 5.2 Semi-supervised Learning
For this setting, we evaluate ELBO, Latent NLL, and Acc. We choose classification accuracy since classification is a common downstream task for semi-supervised learning. We use the same set of priors and baselines. Details on how we partition each prior into $p_{\boldsymbol{{\boldsymbol{\theta}}}}(\mathbf{z}|\mathbf{x},l)$ can be founded in Appendix F. The partitions defined for our priors are local parts of the priors. We note that unlike unsupervised learning, we use the simplified sleep term in our objective for this setting (see Appendix B for details), since $q_{\boldsymbol{\phi}}$ already gets extra information from $l$ here.
The results are shown in Table 2 and Table 8 in Appendix. DDVI mostly outperforms the baselines across different priors and metrics, especially on CIFAR-10 where DDVI is best across the board. For MNIST, DDVI always achieves the best ELBO, and it also performs competitively with other baselines in classification accuracy. We also show the visualizations of the latents in Figure 3 where DDVI matches the prior almost perfectly.
Method Cluster Purity Cluster Completeness NMI
AEVB $0.28\pm 0.02$ $\mathbf{0.78\pm 0.16}$ $0.59\pm 0.08$
AEVB-IAF $0.29\pm 0.04$ $\mathbf{0.73\pm 0.06}$ $0.55\pm 0.06$
AAEB $0.37\pm 0.06$ $\mathbf{0.76\pm 0.11}$ ${0.63\pm 0.02}$
DDVI $\mathbf{0.45\pm 0.03}$ $\mathbf{0.75\pm 0.05}$ $\mathbf{0.66\pm 0.04}$
Table 3: Quantitative genotype clustering results.
@@ 5.3 Clustering and Visualization for Genotype Analysis
In Table 3, we see that DDVI attains the best performance on cluster purity and NMI. For cluster completeness, VAE and AAE have better means but much larger confidence interval. Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA (Wold, Esbensen, and Geladi 1987), t-SNE (Van der Maaten and Hinton 2008), and UMAP (McInnes, Healy, and Melville 2018)). The legend of Figure 4 can be founded at Figure 5 in Appendix.
@@ 6 Discussion
Our approach is most similar to flow-based approximators (Rezende and Mohamed 2015; Kingma et al. 2016); in fact when $T\to\infty$, our diffusion-based posterior effectively becomes a continuous-time normalizing flow (Song, Meng, and Ermon 2020). However, classical flow-based methods require invertible architectures for each flow layer: this constrains their expressivity and requires backpropagating through potentially a very deep network. Our approach, on the other hand, trains a model (a continuous-time flow when $T\to\infty$) via a denoising objective (similar to score matching) that does not require invertible architectures and effectively admits an infinite number of layers (with weight sharing). This model is trained not by backpropagating through the ELBO, but rather via an auxiliary diffusion loss term (effectively, a score matching objective).
@@
Variational posteriors based on GANs (Makhzani et al. 2015) also admit expressive architectures and require only sample-based access to the prior $p(\mathbf{z})$. Our diffusion-based approach admits a more stable loss, and is potentially more expressive, as it effectively supports an infnite number of layers (with shared parameters when $T\to\infty$). Unlike GANs, our models also admit explicit likelihoods and allow us to compute the ELBO for model evaluation. Our approach is similar to variational MCMC (Salimans, Kingma, and Welling 2015); however, we train with a better objective augmented with a diffusion loss, and we adopt improved architectures with shared weights across layers.
@@
Our algorithms are also simpler. For example, diffusion sampling methods require backpropagating through a sampling process to minimize the reverse $\text{KL}(q_{\boldsymbol{\phi}}||p_{\boldsymbol{\theta}})$, which poses challenges with optimization and credit assignment. Some methods based on Schrodinger bridges require an iterative optimization process generalizing the sinkhorn algorithm or computationally expensive on-policy or off-policy (Malkin et al. 2022) trajectory-based optimization. In contrast, DDVI optimizes the forward $\text{KL}(p_{\boldsymbol{\theta}}||q_{\boldsymbol{\phi}})$ using simple gradient-based optimization that directly emulates diffusion-based training.
@@ 7 Related Work
Vahdat, Kreis, and Kautz (2021); Wehenkel and Louppe (2021); Rombach et al. (2022) perform diffusion in the latent space of a VAE. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors.
Our goal is different: we seek to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$. Recent work (Preechakul et al. 2022; Zhang, Zhao, and Lin 2022; Wang et al. 2023) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. Cohen et al. (2022) crafts a diffusion bridge linking a continuously coded vector to a non-informative prior distribution.
@@
Diffusion sampling (Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Zhang et al. 2023; Richter, Berner, and Liu 2023; Sendera et al. 2024; Akhound-Sadegh et al. 2024) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift). Some other works (Zhang et al. 2023; Akhound-Sadegh et al. 2024; Sendera et al. 2024) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. While this work is applicable to our setting, it does not rely on the structure of $p(\mathbf{x},\mathbf{z})$ available to us, namely tractable sampling in $p$.
@@
Latent variable models in general are an attractive alternative to visualization methods like PCA, Sparse PCA, NMF, UMAP, and t-SNE (Wold, Esbensen, and Geladi 1987; Kuleshov 2013; Lee and Seung 2000; McInnes, Healy, and Melville 2018; Van der Maaten and Hinton 2008). Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space. Nevertheless, downsides of LVMs are that they are more computationally expensive and require careful hyperparameter tuning.
@@ 8 Conclusion
While this paper focuses on applications of DDVI to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality. Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making (Nguyen and Grover 2022; Deshpande and Kuleshov 2023), meta-learning (Rastogi et al. 2023), or causal effect estimation (Deshpande et al. 2022).
Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models. Also, attaining competitive likelihood estimation requires architecture improvements that are orthogonal to this paper. However, our ability to generate diverse samples and achieve class separation in latent space hints at the method’s potential on these tasks.
Appendix
@@ Appendix A Pseudocode
Here we provide a pseudocode to illustrate the training process of DDVI.
@@
1:(Optional) Pre-train $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})$ with DDVI but with unconditional diffusion model $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$
@@
4:         $\mathbf{y}_{i}\sim q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}_{i})$ and $\mathbf{z}_{i}\sim q(\mathbf{z}|\mathbf{y}_{i},\mathbf{x}_{i})$ for $i=1,\ldots,k$
5:         Optimize ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ with respect to a Monte Carlo estimate of $\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\|p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z})\right)$ for each $\mathbf{x}_{i}$triangleright Standard ELBO training part
@@
10:              Optimize ${\boldsymbol{\phi}}$ using the standard diffusion noise prediction loss on $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{i},\hat{\mathbf{x}}_{i})$
@@ Appendix B Simplifying Wake-Sleep
In wake-sleep, sampling $\mathbf{x}$ from $p_{\boldsymbol{\theta}}$ to obtain gradients for the sleep term introduces computational overhead. To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ of $\mathcal{L}$:
@@
^$\hat{\mathcal{L}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}\underbrace{-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}\underbrace{-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$}}.$ (12)
@@
We have replaced $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ with a latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$, in which $\mathbf{x}$ is given, and we only seek to fit the true reverse noising process $r(\mathbf{z}|\mathbf{y})$ independently of $\mathbf{x}$. We can similarly show that
$\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$ $=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})}[\log{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\boldsymbol{\theta}})\geq\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})r(\mathbf{y}|\mathbf{z})}[\log({q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}/r(\mathbf{y}|\mathbf{z}))]+\bar{H}(p_{\boldsymbol{\theta}})$ (13)
$=-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})}[\text{KL}(r(\mathbf{y}|\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z})||q(\mathbf{z}|\mathbf{x})),$ (14)
@@
where $\bar{H}(p_{\boldsymbol{\theta}})$ is an entropy term constant in ${\boldsymbol{\phi}}$. Thus, we minimize the forward KL divergence by sampling $\mathbf{z}$, and applying the noising process to get $\mathbf{y}$; the $q_{\boldsymbol{\phi}}$ is fit to denoise $\mathbf{z}$ from $\mathbf{y}$ as in Equation 8.
@@
We optimize our bound on ^$\hat{\mathcal{L}}(x,{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ end-to-end using minibatch gradient descent over ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$. While the wake term is a reconstruction loss as in wake-sleep, the sleep term generates latent samples $\mathbf{z},\mathbf{y}$ from $r(\mathbf{y}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ (by analogy with $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ in normal wake-sleep); the denoiser $q_{\boldsymbol{\phi}}$ is trained to recover $\mathbf{z}$ from $\mathbf{y}$. Thus, we perform wake-sleep in latent space, which obviates the need for alternating wake and sleep phases, and allows efficient end-to-end training. A limitation of this approximation is that the sleep term does not fit $q_{\boldsymbol{\phi}}$ to the true $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x},\mathbf{y})$, and as a consequence $\hat{L}$ is not a tight lower bound on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$. We may think of $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$ as a regularizer to the ELBO.